OpenStudy (hiyathere_20):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1).

OpenStudy (retireed):

Okay, I figured it out, I think. In order to get the roots, you need to use (x+9)(x+3) = 0 which is x^2 + 12x + 27 = 0 That quadratic equation yields a vertex at -6, -9 in order to get the vertex to be at -6, -1. I had to divide the equation by 9. (x+9)(x+3)/9 = 0 Plot it on a graphing calculator.

OpenStudy (hiyathere_20):

okay how do I plot on a graphing calculator

OpenStudy (hiyathere_20):

I got 3 I multiplied 9*3 and got 27 and divided that by 9 and got 3.

OpenStudy (retireed):

demos.com/calculator it looks like this... attached file

OpenStudy (hiyathere_20):

okay so would it be 3x^2